OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 4x – 3y = –6 –x – 5y = –9

OpenStudy (anonymous):

4x – 3(–5y + 9) = –6 4(–5y + 9) – 3y = –6 4x – 3(5y – 9) = –6 4(5y – 9) – 3y = –6

OpenStudy (amistre64):

you have the power ....

OpenStudy (amistre64):

solve it by eliminating the obvious wrong answers

OpenStudy (amistre64):

just keep an eye on the sign of 5 when you solve for the x